### v2

# UniLGL: Learning Uniform Place Recognition for FOV-limited/Panoramic LiDAR Global Localization

Abstract
LiDAR-based Global Localization (LGL) is an essential ingredient for autonomous mobile robots, owing to its robustness against illumination. However, existing LGL methods typically consider only partial information (e.g., geometric features) from LiDAR observations or are designed for homogeneous LiDAR sensors, overlooking the uniformity in LGL. In this work, a uniform LGL method is proposed, termed UniLGL, which simultaneously achieves spatial and material uniformity, as well as sensor-type uniformity. The key idea of the proposed method is to encode the complete point cloud, which contains both geometric and material information, into a pair of Bird’s Eye View (BEV) images (i.e., a spatial BEV image and an intensity BEV image), thereby transforming the LGL problem into a cascaded LiDAR Place Recognition (LPR) and pose estimation problem from the perspective of image fusion. An end-to-end multi-BEV fusion network is designed to extract uniform features, equipping UniLGL with spatial and material uniformity. To ensure robust LGL across heterogeneous LiDAR sensors, a viewpoint invariance hypothesis is introduced, which replaces the conventional translation equivariance assumption commonly used in existing LPR networks and supervises UniLGL to achieve sensor-type uniformity in both global descriptors and local feature representations. Moreover, UniLGL introduces a pipeline that leverages a pre-trained single-image vision foundation model for feature extraction to enhance the multi-BEV fusion LPR network, enabling strong generalization with only a few LiDAR data for fine-tuning. Finally, based on the mapping between local features on the 2D BEV image and the point cloud, a robust global pose estimator is derived that determines the global minimum of the global pose on without requiring additional registration. To validate the effectiveness of the proposed uniform LGL, extensive benchmarks are conducted in real-world environments, and the results show that the proposed UniLGL is demonstratively competitive compared to other State-of-the-Art (SOTA) LGL methods. Furthermore, UniLGL has been deployed on diverse platforms, including full-size trucks and agile Micro Aerial Vehicles (MAVs), to enable high-precision localization and mapping as well as multi-MAV collaborative exploration in port and forest environments, demonstrating the applicability of UniLGL in industrial and field scenarios. The code will be released at https://github.com/shenhm516/UniLGL, and a demonstration video is available at https://youtu.be/p8D-sxq8ygI.
Index Terms:
Global localization, Place recognition, Simultaneous Localization and Mapping (SLAM), Deep learning.I Introduction
Global localization is a fundamental task in developing autonomous robotic systems, which has facilitated various industrial applications, such as autonomous valet parking [1, 2], building inspection [3, 4], and autonomous delivery [5, 6]. Different from ego-motion estimation systems (e.g., visual/LiDAR odometry[7, 8]), which typically assumes the initial pose of the robot to be an identity homogeneous transformation matrix and suffer from drift accumulated by the inaccurate estimation, global localization aims to estimate the global (absolute) pose of robot without prior information. During the last decade, global localization has been achieved using the GNSS[9], infrastructures (e.g., UWB[4], RFIDs[10], and QR codes[11]), and visual/LiDAR place recognition[12, 13, 14, 15, 16]. However, GNSS cannot work in an indoor or dense urban area due to the multi-path effect[9], and infrastructure-based global localization[10, 11, 4] relies on installation and calibration. Besides, visual place recognition can degrade significantly in conditions of appearance changes, i.e., the appearances of particular areas may change drastically under different illumination conditions. Owning to the direct depth measurement of LiDAR, which is immune to scene illumination changes, in recent years, LPR has been widely adopted to enable infrastructure-free global localization in GNSS-denied environments. A typical LGL pipeline consists of two stages: LPR and global pose estimation. Many existing approaches concentrate solely on LPR[14, 17, 18, 19, 20], assuming that global pose estimation can be achieved using point cloud registration algorithms such as ICP[21], GICP[22], or NDT[23]. However, point cloud registration is prone to falling into local minima when there is a large initial pose discrepancy between point clouds, e.g., reverse-direction loop closures. This motivates us to develop a full-fledged LGL system that bridges the gap between topological and metric localization.
Several recent works have been proposed to address the LGL task by leveraging deep learning or exquisitely designed handcrafted descriptors. Although these methods achieve impressive results, they typically utilize only partial information from the point cloud or are tailored to specific types of LiDAR sensors. LiDAR inherently captures both spatial structure and material properties, represented in 4D data comprising 3D coordinates and intensity. However, methods such as [24, 15, 25, 19, 16, 26, 27] either neglect the material cues encoded in the intensity channel or exploit only partial spatial information, treating LiDAR as a 3D sensor. As an example shown in Fig. 1(a), if only the 3D spatial information of the point cloud is considered, high-reflectivity painted text on the ground will be neglected. This pruning of information can compromise the performance of global localization. Moreover, most existing learning-based LGL methods[24, 15, 25] follow the paradigm of position-based place recognition [28], which assesses whether a robot revisits the same geographical region despite changes in viewpoint. However, as illustrated in Fig. 1(b), for Field of View (FoV)-limited LiDARs, viewpoint invariance is not equivalent to rotation invariance. In such sensors, the observed point clouds at the same position but with different orientations can be almost disjoint, making conventional rotation-invariant designs insufficient for achieving viewpoint-invariant global localization. In view of the aforementioned analysis, a question arises: Is it possible to learn uniform place recognition that enables LGL preserving spatial and material uniformity as well as sensor-type uniformity?
I-A Related Works
I-A1 LPR
In early works [29, 30, 31] of LPR, point statistics (such as signature and histogram) are commonly exploited to represent the point cloud appearance. M2DP [29] projects a point cloud to multiple 2D planes and generates a density signature for points in each plane. The Singular Value Decomposition (SVD) components of the signature are then used to compute a global descriptor. Under the assumption that the point cloud subject to a Gaussian Mixture Model (GMM), the 3D-NDT histogram [30] constructs a place similarity metric for LPR based on the number of linear, planar, and spherical distributions within the GMM, where Principal Component Analysis (PCA) is applied to each Gaussian component to classify it into one of the three categories including linear, planar, and spherical. Authors of [31] extend the 3D-NDT histogram [30] to enable fast LPR for FoV-limited LiDARs and integrate it with a LiDAR odometry to build a complete SLAM framework, called LOAM-Livox [32], which is tailored for FoV-limited LiDAR sensors. However, since the histogram-based method only provides a stochastic index of the scene, it fails to explicitly capture detailed structural information, which limitation reduces the descriptor’s discriminability for place recognition.
To overcome this limitation, Scan Context [19] is proposed, which divides the 3D space into bins in a polar coordinate and records the maximum height of the points in each bin. In this way, the 3D spatial structure of a point cloud is encoded into a compact 2D image representation. Place recognition is then achieved by measuring the similarity between the images generated from different scans. As an extension of Scan Context [19], Scan Context++ [16] proposes a novel spatial division strategy to improve lateral invariance and integrates place retrieval with coarse yaw alignment as the initial guess of point cloud registration. While many handcrafted approaches have been developed for extracting informative LPR descriptors, inspired by the remarkable success of deep networks in natural language processing and computer vision, discriminative learning-based LPR methods have emerged and demonstrated competitive performance. PointNetVLAD [14] established a foundational LPR framework by extracting local features from raw point clouds using PointNet [33] and aggregating local features into a global descriptor via a NetVLAD[12, 13] pooling layer. The resulting global descriptor is then used for LPR through K-Nearest Neighbor (KNN) search. MinkLoc3D [34] employs sparse 3D convolutions for local feature extraction and introduces a Generalized-Mean (GeM) pooling layer to aggregate the local features into a global descriptor. HeLioS [18] employs a U-Net-style architecture with sparse convolution[35] to extract local features from a point cloud, which are then aggregated into LPR global descriptors through GeM pooling and the SALAD[36] attention module. LoGG3D-Net[37] introduces a local consistency loss to guide the LPR network in learning consistent local features across revisits, and high-order aggregation has also been explored to enhance the repeatability of global descriptors, which results in an overall improvement in LPR performance. However, these methods all rely on point cloud input, which is characterized by their sparsity and lack of orderly structure. Some methods explore the potential of image-based representations instead of direct point-based representations. OverlapNet [38] represents point clouds in Range Image View (RIV) and employs a Siamese network to estimate the overlap between two range images, enabling robust and generalizable LPR. OverlapTransformer [39] builds upon OverlapNet [38] by incorporating a transformer-based attention mechanism [40] to enhance performance. ImLPR [17] leverages the vision transformer foundation model, DINOv2 [41], to further boost LPR performance through RIV representations.
I-A2 LGL
For the aforementioned methods, LiDAR measurements are solely used for place recognition, under the assumption that the relative pose between the corresponding point clouds can be estimated using point cloud registration techniques. However, point cloud registration is a typical non-convex optimization problem, and conventional registration methods[21, 22, 23] rely on a rough initial guess to ensure that the relative pose converges to the global minimal. Although some approaches [42, 43] attempt to overcome this limitation via convex relaxation, they often require additional point cloud feature extraction and matching. To address this issue, a series of works attempt to design a complete LGL system that performs both place recognition and relative pose estimation in a unified framework. In early works on LGL, local feature extractors directly applied to 3D point clouds (e.g., SHOT[44], FPFH[45], and 3D SIFT[46]) as well as those applied to image-represented point clouds(e.g., ORB[47] and Harris[48]) were utilized to achieve LPR and relative pose estimation through local feature matching. These local-feature-based LGL methods demonstrate promising performance in small-scale environments but often lack sufficient distinctiveness for large-scale outdoor scenes.
To address this limitation, recent research further aggregates global descriptors from local features, enhancing robustness to local noise and varying point cloud densities, while the local features are used for corresponding point cloud alignment. BoW3D [49] extracts local features through Link3D [50] and adopts Bag of Words[51], which is widely adopted in visual place recognition fields, as the global descriptor for LPR. The relative pose estimation is achieved by local feature matching and alignment. STD [26] voxelizes the point cloud to extract keypoints as local features and constructs a triangle descriptor based on voxel distribution for place recognition. The relative pose between matched point clouds is estimated by aligning the triangle vertices using the Umeyama alignment [21]. BTC [27] improves global localization performance by introducing a binary descriptor into the STD[26], providing a more detailed and discriminative representation of local point cloud geometry. RING [52] and RING++ [53] apply the Radon transform to the BEV image of point clouds to extract local features. A translation-invariant global descriptor is then constructed by applying the discrete Fourier transform to these local features. To achieve both LPR and relative pose estimation, an exhaustive search over the orientation space is performed to maximize the circular cross-correlation between the query and database global descriptors. Subsequently, the relative translation between the corresponding point clouds can be solved in closed form over the frequency domain using the local feature information. The handcrafted descriptors used in the aforementioned methods, though elegantly designed, often suffer from low information density. To mitigate the risk of numerous false matches, handcrafted LPR is typically coupled with a geometric verification step. For example, STD [26] and BTC[27] typically identify multiple-loop candidate point clouds for a query scan through descriptor voting, and then select the one with the best geometric verification. Similarly, RING [52] and RING++ [53] directly couple rotation estimation with the LPR process by performing an exhaustive search over the orientation space.
To address this challenge, a series of methods have been proposed to learn highly discriminative descriptors. LCDNet [54] presents an end-to-end LGL framework, which employs PV-RCNN [55] for robust local feature extraction and NetVLAD [12, 13] for global descriptor aggregation. LCRNet [56] introduces a novel feature extraction backbone and a pose-aware attention mechanism to jointly estimate place similarity and 6-DoF relative pose between pairs of LiDAR scans. SpectralGV [57] extends Logg3D-Net[37] to LGL by leveraging its local features to register point cloud pairs that are initially matched using the global descriptors produced by Logg3D-Net[37]. BEV-Place[24] and BEV-Place++[15] represent point clouds using BEV images and adopt a rotation equivariant and invariant network to extract local features from the BEV images. The local features are then aggregated into a global descriptor using NetVLAD[12, 13]. BEV-Place[24] and BEV-Place++[15] demonstrate that representing point clouds with BEV images yields superior generalization capability compared to raw point clouds, particularly under viewpoint variations and scene changes. Similarly, RING# [25] proposes a feature learning architecture to simultaneously learn LPR and 3-DoF global localization, leveraging the frequency-domain BEV representation introduced in RING[52] and RING++[53].
I-A3 LPR/LGL with Uniformity
Despite the extensive research on the fundamental techniques underlying LPR and LGL, the concept of uniformity has rarely been considered. To achieve spatial and material uniformity, a series of methods integrate the intensity information into well-established place recognition descriptors. In [58], a uniform LPR descriptor is proposed, which encodes intensity information into the original SHOT descriptor[44] constructed from 3D geometric information. Intensity Scan Context [20] replaces the height channel used in the original Scan Context [19] with the intensity of LiDAR points, which justifies that intensity information can be distinctive for places recognition. In[59], intensity readings are utilized to generate an intensity image, which serves as an alternative to the commonly used RIV image in conventional LPR pipelines. In addition to the aforementioned handcrafted descriptor-based LPR or LGL methods, some approaches, such as LoGG3D-Net[37] and LCDNet[54], directly take 4D point clouds (containing both geometric and intensity information) as network input, attempting to learn spatial and material uniformity. Considering the sparsity of point clouds, methods like OverlapNet[38] and ImLPR[17] represent the 4D point cloud using multi-channel RIV images including range channel, intensity channel, and normal channel, etc., and learn the similarity between a pair of images to achieve spatial and material uniform LPR.
Compared to spatial and material uniformity, sensor-type uniformity has received attention in only a few handcrafted descriptor-based LPR methods [26, 27, 52, 53]. As corroborated by Solid [60], designing an LPR system with sensor-type uniformity can significantly improve performance, particularly for FoV-limited LiDARs. However, most existing learning-based LPR/LGL methods [24, 15, 25, 54, 56, 37, 33, 17] follow the paradigm of position-based place recognition [28], where networks are supervised based on the geographical distance between scans to determine whether a robot revisits the same place. These approaches implicitly assume that LiDAR observations are rotation-invariant. As illustrated in Fig. 1(b), panoramic LiDARs naturally satisfy this assumption due to their omnidirectional sensing capability. In contrast, FoV-limited LiDARs’ observations at the same location under different headings may be completely disjoint. OverlapNet [38], OverlapTransformer[39], and HeLioS[18] offer an alternative by supervising the network using overlap rather than geographical distance, thus providing a more natural metric than geographical distance for point cloud similarity. However, HeLioS[18] only considers the sensor-type uniformity at the global descriptor level by using point clouds as input to the LPR network, which neglects such uniformity at the local feature level, restricts it to LPR without metric localization. OverlapNet [38] and OverlapTransformer[39] represent LiDAR scans as RIV images, which limits their generalization to FoV-limited LiDARs. Specifically, to ensure that range images do not contain large regions of vacant pixels, different horizontal resolutions have to be used for panoramic and FoV-limited LiDARs, making it difficult to generalize a single model across heterogeneous LiDAR configurations.
Capability |
Place Recognition | Handcrafted | Handcrafted | Handcrafted | Learning | Learning | Learning | Learning | Learning | Learning | Learning | Learning | ||||||||||||||||||||||
[…]
Point Cloud Representation | Point | Point | BEV | Point | RIV | RIV | BEV | BEV | Point | Point | BEV |
I-B Motivation and Contributions
In view of the aforementioned analysis, as summarized in Table I, most existing methods either focus exclusively on LPR[17, 20, 18] or provide only partial solutions for LGL, such as 1-DoF heading alignment[60, 38] or 3-DoF pose estimation constrained to SE(2)[15, 53, 25]. Moreover, the uniformity illustrated in Fig. 1 is commonly neglected [15, 25] or only partially considered [27, 20, 60, 38, 54, 37, 17, 18]. These limitations motivate us to develop a uniform LGL system, called UniLGL, that enables fully 6-DoF global pose estimation over SE(3), while simultaneously preserving spatial and material uniformity as well as sensor-type uniformity. The key idea of UniLGL is to represent the 4D point cloud using two BEV images that encode both geometric and intensity information. A novel feature fusion network is then designed elaborately to learn local features and global descriptors from the BEV images, supervised under the viewpoint invariance hypothesis. Considering the strong cross-task generalization capabilities of foundation models, as illustrated by the t-SNE[61] visualization in Fig. 2, initializing the LPR network with a foundation model (without fine-tuning) endows the LPR network with an initial capability for place recognition. As shown in Table I, foundation model has rarely been explored in the context of LPR tasks. This motivates us to explore leveraging foundation models to empower the LPR network, and to bridge the domain gap between foundation models and the LPR task through fine-tuning with a small amount of LiDAR data, enabling high-performance place recognition, as illustrated in Fig. 2(c). Subsequently, uniform LGL is realized by combining uniform LPR using global descriptor matching with 6-DoF global pose recovery through local feature matching. To recap, the main contributions of this paper are listed as follows:
Uniform LPR: An end-to-end LPR network is designed to provide a uniform place representation. UniLGL fuses the spatial BEV images and intensity BEV images of LiDAR scans through a novel feature fusion network to achieve spatial and material uniformity, and a viewpoint invariance hypothesis is introduced to supervise UniLGL with sensor-type uniformity, which hypothesis replaces the conventional translation equivariance hypothesis commonly used in conventional LPR networks.
Global Localization: UniLGL provides a complete global localization framework that achieves both LPR and 6-DoF pose estimation on SE(3) without requiring additional point cloud registration. Unlike conventional image-based (BEV, RIV, etc.) LGL methods, which are typically limited to 1-DoF or 3-DoF pose estimation, the proposed method enforces local feature consistency within the network. This enables robust 6-DoF global localization through local feature matching between BEV images.
Foundation Model: Following the paradigm shift of the task-agnostic foundation model in Natural Language Processing (NLP) and computer vision, we explore the incorporation of foundation models into LGL, and an adaptation strategy is proposed to incorporate the foundation model originally designed for single-image feature extraction into our multi-BEV-image fusion network. By leveraging the strong generalization capability of foundation models, UniLGL is able to deliver effective performance using only a small amount of LiDAR data for fine-tuning.
Fully-fledged: UniLGL is a full-fledged LGL framework that has been validated across multiple public datasets as well as real-world applications. Extensive benchmark comparisons on public datasets demonstrate that UniLGL delivers competitive place recognition and global localization performance compared with SOTA LGL methods. Beyond benchmark evaluations, UniLGL has been further extended to various real-world industrial and field applications, including autonomous driving trucks and collaborative exploration with multi-MAV systems, to demonstrate its performance and industrial applicability.
II Learning Uniform Global Descriptor for LiDAR Place Recognition
Given a query point cloud observation and a set of database point cloud , LPR is in charge of retrieving the most similar point cloud to from .
where measures the similarity of two point clouds. In this work, a mapping function is developed that represents the point cloud with a global descriptor and transforms the LPR problem (1) into a descriptor matching problem.
where and are descriptors of and , respectively. The descriptor matching problem (2) can be easily solved by performing a KNN search. Therefore, the key to solving the LPR problem (1) lies in learning a mapping function that ensures point clouds with a high similarity yield highly similar descriptors, and vice versa.
Remark 1
(Uniform LPR): In this work, the term Uniform manifests in two key aspects: 1) Spatial and Material Uniformity: the proposed UniLGL leverages both the spatial structure and material information (i.e., intensity) captured by LiDAR for place recognition; and 2) Sensor-Type Uniformity: it is appropriate for both FoV-limited and panoramic LiDAR.
II-A Learning Spatial and Material Uniformity
LiDAR is a 4D sensor that captures point clouds , where each point is represented by its 3D spatial coordinates and intensity . However, most existing LPR methods [24, 15, 25, 19, 16, 26, 27, 20, 59] overlook intensity or partially discard geometric information. To address this limitation, we propose representing the point cloud using two complementary BEV images, called spatial BEV image and intensity BEV image , that jointly preserve both spatial and intensity information.
where denotes the pixel value of the pixel , denotes the point cloud corresponding to the pixel , is the resolution of BEV images, denotes the floor operation, returns the cardinal number of a set, and returns the intensity values of a point cloud.
To extract the uniform descriptor of the spatial and intensity BEV images, we extend the impressive ViT [62] network to enable feature fusion. As illustrated in Fig. 3, BEV images are split into a sequence of flattened 2D patches , where is the resolution of each image patch. Two independent Convolutional Neural Networks (CNNs) are employed to learn the patch embedding projection for spatial and intensity BEV images, respectively, and a learnable classification vector is augmented to the sequence of embedded patches.
where and denote the -th patch of the spatial BEV image and the intensity BEV image, respectively; and represent the learnable embedding projections for the spatial and intensity BEV images, and is the dimension of each embedded patch. The resulting embedded patches are concatenated with the classification vector CLS into , where is the number of patches per BEV image. A standard learnable position embedding is employed on to retain positional information, the resulting sequence of embedding vectors serves as input to the transformer encoder.
where , , and are position embeddings of classification vector CLS, embedded patches of spatial BEV image, and embedded patches of intensity BEV image, respectively.
The Transformer encoder[63], composed of alternating Multi-head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) blocks, is employed to learn both the global descriptor (i.e., classification token) for LPR and the local features (i.e., local feature token) of each patch.
where is the joint-image-level classification token, and and are patch-level local feature tokens, respectively. The joint-image-level classification token is treated as the unified global descriptor for LPR.
Remark 2
(Spatial and Material Uniformity) For now, most existing LPR methods [24, 15, 25, 19, 16, 26, 27, 18] treat LiDAR solely as a 3D geometric sensor, neglecting the material properties encoded in the intensity channel. It has been corroborated in [20, 59] that incorporating intensity information into LPR can significantly improve performance. However, although [20] and [59] incorporate both geometric and intensity cues into BEV and intensity RIV images, respectively, the projection to 2D image inevitably causes a loss of 3D geometric fidelity, as [20] retains only and [59] retains only . To mitigate the information loss, we separately preserve geometric and intensity cues using a spatial BEV image and an intensity BEV image. Built upon the single-image feature extraction backbone, ViT[62], we design a novel feature fusion network that enables joint processing of multiple images, facilitating uniform representation learning across both spatial (geometry) and material (intensity) domains.
II-B Learning Sensor-Type Uniformity
In this section, we begin a discussion with the Translation Equivariance hypothesis, which is a fundamental requirement in conventional LPR[24, 15, 14, 54, 25].
Hypothesis 1
(Translation Equivariance) An effective global descriptor should ensure that point clouds captured at spatially proximate states yield highly similar descriptors, and vice versa.
where denotes the translation state of the robot capturing the query point cloud, and and denote the translation states of the robot capturing the database point cloud.
Definition 1
(Rotation Invariance) If a mapping function is rotation invariant, it satisfies the following equation.
where denotes the point cloud transformed by an arbitrary rotation .
An important implication of Hypothesis 1 is that the global descriptor extraction network should exhibit rotation invariance, ensuring that point clouds captured at the same location but under different rotations yield identical descriptors. However, LPR networks with built-in rotation invariance are not well-suited for FoV-limited LiDARs. As illustrated in Fig. 1(b), panoramic LiDARs capture highly overlapping point cloud data at nearby locations even under different orientations, whereas FoV-limited LiDARs may observe almost disjoint point clouds due to their restricted FOV. To mitigate this limitation, a Viewpoint Invariance hypothesis is proposed to supervise the learning of the global descriptor extraction network, encouraging it to produce consistent descriptors under different viewpoints. As shown in Fig. 4, viewpoint invariance serves as a unified hypothesis for both FoV-limited and panoramic LiDAR, which defines positive point clouds based on scene similarity rather than spatial proximity. Moreover, for panoramic LiDARs, LPR methods based on the Viewpoint Invariance hypothesis are expected to achieve superior performance compared to those relying on the Translation Equivariance hypothesis. For example, translation-equivariance-based LPR methods [24, 15, 14, 54, 25] typically use a distance-based criterion and would treat point clouds with substantial co-visible regions (as shown in Fig.4) as negative samples due to their spatial separation. In contrast, under the viewpoint invariance hypothesis, the LPR network not only recognizes spatially close point clouds but can also identify distant point clouds with co-visible areas.
Hypothesis 2
(Viewpoint Invariance) An effective global descriptor extraction network should ensure that point clouds capturing similar scenes from arbitrary viewpoints yield similar descriptors, and conversely, that dissimilar scenes produce dissimilar descriptors.
Hypothesis 2 defines viewpoint invariance as a quantitative hypothesis. To supervise the global feature extraction network accordingly, a mathematical formulation of viewpoint invariance, similar to the conventional translation equivariance defined in (7), is required. The Intersection over Union (IoU) of the convex hulls of point clouds is introduced to determine whether they share substantial co-visible regions.
where returns the convex hull of a point cloud, denotes the area measurement, and and represent the query and -th database point cloud, respectively.
According to the mathematics formulation (9) of the substantial co-visible region of point clouds, Viewpoint Invariance can be formulated as
An intuitive approach to enforcing sensor-type uniformity in LPR is to supervise the descriptor extraction network using the IoU (9) between the convex hull of the query point cloud and those of the positive and negative samples, respectively. The lazy triplet loss[14] is adopted to supervise the place recognition process, which aims to maximize the descriptor distance between a query and negative point clouds while minimizing the descriptor distance between a query and a positive image.
where is the LPR loss, is a constant margin, denotes the negative global descriptor set, and is the positive global descriptor that corresponds to the query descriptor . For each query point cloud, its positive samples are convex hulls that satisfy IoU , and the negative samples are images with IoU .
Given that UniLGL is proposed to jointly achieve LPR and 6-DoF pose estimation, the local feature tokens have to be explicitly supervised for viewpoint invariance. For each image pair matched via LPR, we divide both images into multiple patches and randomly select one pixel from each patch as a keypoint. A keypoint from located in the overlapping region between and can find its corresponding keypoint in using the ground truth trajectory. The pixel is treated as the positive pixel of while all other keypoints on are considered negative samples. To enable pixel-level contrastive learning, each pixel of an image is represented by a local feature , obtained via interpolation from the patch-level local feature token map. We introduce the InfoNCE loss[64] to maximize the cosine similarity between the local feature and its corresponding feature while minimizing the cosine similarity between and its negative samples.
where denotes the local feature loss of a matched image pair , which is designed to supervise the viewpoint invariance in local feature perspective; is the set of local features in located within the overlapping region between and ; and is the full local feature set of .
To achieve sensor-type uniformity in both LPR and global pose estimation perspective, the final loss function is designed as a linear combination of the LPR loss (11) and local feature loss (12):
where is the loss balancing coefficient, empirically set to , and and denote the spatial BEV image pair and the intensity BEV image pair, respectively.
II-C LPR Meets Foundation Models
Motivated by the remarkable success of task-agnostic pre-trained representations [65, 66, 67] in NLP, we explore the incorporation of foundation models into the proposed UniLGL, which are trained on large-scale datasets to learn general-purpose features. As noted in Remark 3, owning to the efficiency for acquiring large-scale visual data, in this paper, a self-supervised visual feature extraction foundation model, DINO vision transformer[68], is adopted to pre-initialize the feature fusion network designed in Section II-A.
where and represent the embedding and positional encoding layers, respectively, and and denote the MSA and MLP networks, all pre-trained as part of the DINO vision transformer. In (14), we expand the use of DINO pre-trained weights, initially developed for single-image feature extraction, to support the proposed multi-image feature fusion LPR architecture.
Remark 3
(Why LPR Meets Foundation Models) A hypothesis that has been corroborated in [69, 70, 12, 13, 71] is that the high performance of most modern place recognition can be mainly attributed to large-scale training. Over the past decade, the scale of training data used for place recognition has grown explosively, from hundreds of thousands in early works (e.g. Pittsburgh Dataset [72] powered NetVLAD[12, 13]) to millions (e.g. Google-Landmark datasets[73] powered DeLF[69] and DeLG[70]) or even tens of millions (e.g. San Francisco XL dataset powered CosPlace[71]) in recent years. However, unlike visual images that can be easily acquired at scale through map services (e.g., Google Street View), collecting large-scale point cloud data remains challenging and resource-intensive. Consequently, most existing LPR methods either rely on classical hand-crafted features[16, 27, 19, 52, 53] or lightweight models[24, 15, 25, 54] trained on relatively small-scale datasets, typically comprising only tens of thousands of samples. This motivates the introduction of foundation models pre-trained on large-scale visual data into LPR, achieving high performance through the use of only a small amount of LiDAR data for fine-tuning.
III Robust Global Registration on Manifolds
Global registration is a fundamental task that aims to align a query point cloud with a database point cloud , where is retrieved via a KNN search over the global descriptor space (2). Conventional point cloud registration methods, such as ICP[21], GICP[22], and NDT[23], estimate the relative pose between LiDAR point clouds iteratively with an initial guess. However, in real-world scenarios, many challenging tasks have to be tackled without any prior knowledge, such as life-long navigation, and multi-agent collaborative localization and mapping. As demonstrated in Section II-A, the proposed uniform LPR network extracts not only the classification token, which serves as the global descriptor of a point cloud, but also patch-level local feature tokens, denoted as and . Specifically, the -th patch of the spatial and intensity BEV images is represented by and , respectively, where . To enable pixel-level matching, a dense feature map is reconstructed by interpolating the patch-level local feature tokens, allowing each pixel in the BEV image to be associated with a local feature . Subsequently, for each keypoint in the query image, a correspondence is established by performing a KNN search in the pixel-level local feature embedding space of the database image.
where and denote the -th and -th pixel-level local features of the spatial BEV images from the query and database, respectively. Similarly, and represent the -th and -th pixel-level local features of the intensity BEV images from the query and database, respectively.
Benefiting from the BEV image representation of 3D point clouds (3) and the local feature supervision strategy designed in Section II-B, UniLGL is able to achieve point-level global matching through pixel-level matching (15) on BEV images.
where and denote point clouds corresponding to the -th pixel of the query spatial BEV image and the -th pixel of the database spatial BEV image, respectively. Similarly, and respectively denote point clouds corresponding to the -th pixel query intensity BEV image and the -th pixel of the database intensity BEV image. According to the point-level matching result (16), the global registration problem is defined as:
where and are -th and -th points of the query point cloud , respectively; and are the corresponding points of and in the database point cloud , respectively; and are the number of matched keypoints corresponding to spatial BEV image and intensity BEV image, respectively; is the relative pose between query and database point clouds.
With the consideration of matching outlier, we assume the noise of point matching is unknown but bounded, and write the relative pose estimation problem in a Truncated Least Squares (TLS) formulation:
where and are the -th and -th rows of and , respectively. The TLS formulation of the global registration problem (18) discards measurements with large residuals (when the -th summand does not influence the optimization). To solve the global optimization problem (18) without an initial guess, a graduated non-convexity [74] optimization algorithm is derived. According to the Black-Rangarajan duality [75], the TLS problem (18) is equivalent to:
where and are slack variables. The graduated non-convexity TLS problem (19) can be solved by alternating optimization:
where (20) is a weighted version of the outlier-free pose estimation problem (17), which can be solved globally using SVD, and (21) can also typically be solved in closed form.
IV Experimental Setup
During experiments, the dimension of the joint-image-level classification token and patch-level local feature token is set to , and DINO-ViTs-8[68] is introduced as the foundation model of the proposed UniLGL. We use the Adam optimizer with a learning rate of and a weight decay of . Each experiment conducted in Section V is evaluated on a computer equipped with an Intel Core i9-13900KF and an NVIDIA GeForce GTX 4080. For a fair comparison, all methods listed in Section IV-B are retrained for 20 epochs following their default configurations on sequences ntu_day_01, ntu_night_08, Snail_81R_01, and Garden_db. The details of the data sequences are summarized in Section IV-A.
IV-A Datasets
UniLGL aims to enable uniform global localization for various types of LiDARs in diversified environments. To demonstrate the effectiveness of the proposed method, experiments are conducted on three representative datasets, which encompass a variety of environments including campus areas, urban roadways, and highly repetitive outdoor unstructured scenes.
IV-A1 MCD[76]
The MCD dataset includes both FoV-limited (Livox Mid-70) and panoramic (Ouster OS1-128) LiDAR measurements collected in campus environments. For evaluation, five high-speed sequences, including ntu_day_01, ntu_day_02, ntu_day_10, ntu_night_08, and ntu_night_13, are selected, that feature numerous loop closures and span both daytime and nighttime conditions. During the experiments, ntu_night_08 served as the database sequence. To facilitate a fair comparison of benchmark methods under different LiDAR configurations, we employ the Livox Mid-70 (FoV-limited) and the Ouster OS1-128 (panoramic) as sensor inputs. The associated data sequences are referred to as Mid_NTU_XX and OS_NTU_XX, respectively.
IV-A2 Snail[77]
The Snail dataset was collected using a roof-mounted panoramic LiDAR (Hesai Pandar XT32) and contains extensive urban driving data featuring dynamic objects and high-rise buildings. For evaluation, three sequences, including 20240116-2, 20240123-2, and 20240116-3, are selected, covering a total distance of real urban driving scene over . In the experiments, each of the above three sequences is divided based on travel distance. Specifically, the first of each trajectory is used as the database, and the remaining is used as the query set for evaluating the benchmark methods. The resulting split sequences are named Snail_81R_01, Snail_81R_02, and Snail_81R_03, respectively.
IV-A3 Garden[78]
The Garden dataset was collected with a Husky mobile robot platform mounted with a panoramic LiDAR (Ouster OS1-32) for robot global localization in unstructured and highly repetitive environments, which are characterized by dense vegetation and symmetric, repetitive paths. To evaluate the long-term global localization performance of UniLGL on FoV-limited LiDARs, we further augmented the Garden dataset using an Agile Hunter autonomous mobile robot platform equipped with a FoV-limited LiDAR (Livox Avia) and a panoramic LiDAR (Robosense Helios 32). Five new sequences, namely Garden_db, Garden_01, Garden_02, Garden_03, and Garden_04, are augmented to the original Garden dataset, which are collected in the same region of the Garden dataset. Notably, the original four Garden sequences collected using a Husky robot, referred to as Garden_LT_01, Garden_LT_02, Garden_LT_03, and Garden_LT_04, provide long-term measurements across 8 months when compared with the 5 augmented data sequences. During the experiments, Garden_db is used as the database, while Garden_01 to Garden_04 are employed to evaluate the localization performance of UniLGL on FoV-limited LiDARs. In addition, Garden_LT_01 to Garden_LT_04 are utilized to assess the long-term generalization ability of UniLGL. The 6-DoF ground truth trajectory is obtained by fusing the RTK-GNSS with multi-LiDAR odometry[8] through offline pose graph optimization.
IV-B Comparison Baseline
UniLGL aims to achieve uniform global localization across spatial and material domains, as well as between FoV-limited and panoramic LiDAR observations. To demonstrate the effectiveness of the proposed method, we present detailed quantitative analyses comparing UniLGL with SOTA LPR/LGL methods, including BEVPlace++[15], LoGG3D-Net[37], and RING++[53]. Moreover, to investigate the impact of introducing spatial and material uniformity as well as sensor-type uniformity into LGL, we conduct an ablation study on the proposed UniLGL by evaluating it under various configurations. Details of the baselines are given as follows:
BEV-Place++[15]: a LGL method based on spatial BEV image representation. The BEV-image-represented point cloud enables LPR and global pose recovery through a rotation-invariant image similarity detection network and image registration, respectively.
LoGG3D-Net[37]: an end-to-end LPR method that accounts for spatial and material uniformity by learning global place recognition descriptors directly from raw 4D point clouds, which are composed of 3D spatial coordinates and intensity information. To complement the limitations of LoGG3D-Net as a pure LPR method, SpectralGV[57] is integrated to provide 6-DoF metric localization capability.
RING++[53]: a handcrafted LGL method with sensor-type uniformity, which constructs translation-invariant descriptors and orientation-invariant metrics over the BEV Radon sinogram by leveraging the spatial information of the point cloud.
UniLGL w/o Intensity: utilizes only the spatial information of the point cloud for global localization by extracting LPR descriptors and local features from the spatial BEV image without ensuring spatial and material uniformity.
UniLGL w/o Spatial: utilizes only partial spatial and intensity information from the point cloud for global localization by extracting LPR descriptors and local features from the intensity BEV image while eliminating height information from the point cloud.
UniLGL: the full algorithm proposed in this paper that comprehensively considers spatial and material uniformity together with sensor-type uniformity.
It is worth noting that the aforementioned methods achieve sensor-type uniformity by supervising the viewpoint invariance defined in Hypothesis 2. In contrast, the original BEVPlace++[15] and LoGG3D-Net[37] are designed based on the translation equivariance hypothesis defined in (7). To validate the effectiveness of the proposed viewpoint invariance supervision, we additionally train UniLGL (called UniLGL Dis. Sup.) and the SOTA LGL method BEVPlace++ (called BEVPlace++ Dis. Sup.) using the conventional translation equivariance hypothesis for comparison.
UniLGL Dis. Sup.: UniLGL is supervised by the translation equivariance hypothesis. Following the training strategy used in BEVPlace++, for each query frame, its positive samples are the ones within away from itself and its negative samples are the other frames.
BEVPlace++ Dis. Sup.[15]: BEVPlace++ is supervised by the translation equivariance hypothesis, which is consistent with its original configuration.
V Experimental Evaluations and Validations
V-A Evaluation of Place Recognition
For place recognition evaluation, the IoU/distance between two point clouds is used to determine whether a retrieved match is correct. For methods supervised by the translation equivariance hypothesis (Hypothesis 1), such as UniLGL Dis. Sup. and BEVPlace++ Dis. Sup., point cloud pairs with distances below are chosen as positive place recognition samples. For other methods listed in Section IV-B, point cloud pairs with are chosen as positive place recognition samples. In this experiment, three metrics are leveraged to assess the performance of all methods listed in Section IV-B.
Top-1 Recall: For each query, we find its nearest descriptor and retrieve the Top-1 match from the database. According to the IoU/distance threshold, we determine whether the prediction is a True Positive (TP), False Positive (FP), or False Negative (FN). The Top-1 recall rate is defined as the ratio of TP overall positives:
(22) -
Average Precision: Precision is computed as the ratio of TP overall predicted positives:
(23) By setting different descriptor distance thresholds, the corresponding precision and recall pair can be calculated. The average precision is the area under the Precision-Recall curve.
Precision–recall curve: A curve that plots the precision and recall of the retrieval results as the descriptor distance threshold changes.
Sequence | UniLGL |
BEVPlace++ | Logg3D-Net | RING++ |
FoV-limited LiDAR |
Mid_NTU_02 | 99.87/88.88 | 99.73/85.82 | 98.72/87.35 | 99.20/86.69 | 88.56/60.56 | 80.38/45.22 | 41.63/8.33 | 80.73/75.20 | 59.94/40.34 | ||||||||||
Mid_NTU_10 | 95.75/86.43 | 95.82/83.87 | 95.70/83.95 | 94.60/83.49 | 94.07/60.02 | 91.59/45.04 | 60.98/9.56 | 91.78/79.44 | 92.98/64.29 | |||||||||||
Mid_NTU_13 | 98.71/88.38 | 98.58/85.39 | 98.36/87.21 | 96.71/85.42 | 93.25/64.53 | 76.37/40.62 | 42.03/6.76 | 81.24/74.14 | 56.09/44.23 | |||||||||||
Garden_01 | 97.67/83.92 | 96.81/82.61 | 97.30/83.05 | 95.15/79.77 | 95.24/59.76 | 90.67/54.18 | 86.59/15.60 | 84.33/78.46 | 84.21/35.58 | |||||||||||
Garden_02 | 98.65/84.87 | 98.52/83.12 | 98.19/82.88 | 93.87/78.09 | 94.64/63.19 | 91.02/45.84 | 86.89/16.08 | 81.59/75.70 | 81.20/37.16 | |||||||||||
Garden_03 | 98.63/85.45 | 96.15/82.87 | 97.33/77.54 | 91.57/70.55 | 91.14/56.48 | 84.92/36.64 | 72.77/15.44 | 81.57/73.76 | 82.93/35.27 | |||||||||||
Garden_04 | 99.42/86.61 | 99.18/86.35 | 98.66/83.25 | 96.14/79.77 | 99.17/59.93 | 91.05/41.76 | 81.05/19.53 | 84.88/73.34 | 87.98/32.39 | |||||||||||
Average | 98.38/86.36 | 97.82/84.29 | 97.75/83.60 | 95.32/80.54 | 93.72/60.64 | 86.57/44.19 | 67.42/13.04 | 83.73/75.72 | 80.46/41.32 | |||||||||||
Panoramic LiDAR |
OS_NTU_02 | 100.0/89.93 | 100.0/86.19 | 100.0/87.57 | 100.0/92.83 | 100.0/76.56 | 98.32/74.75 | 77.72/24.52 | 98.48/94.37 | 98.33/94.09 | ||||||||||
OS_NTU_10 | 99.94/88.49 | 99.75/88.95 | 100.0/88.39 | 99.92/90.64 | 99.69/76.76 | 99.10/62.91 | 78.46/26.32 | 100.0/97.23 | 100.0/97.31 | |||||||||||
OS_NTU_13 | 100.0/90.81 | 100.0/87.37 | 100.0/89.23 | 100.0/90.72 | 99.78/76.99 | 88.99/62.77 | 62.28/18.08 | 99.71/92.65 | 99.31/93.09 | |||||||||||
Snail_81R_02 | 95.62/79.79 | 95.03/76.55 | 92.73/71.82 | 96.15/88.76 | 92.08/59.95 | 72.37/30.19 | 36.09/6.03 | 98.67/89.93 | 99.50/73.61 | |||||||||||
Snail_81R_03 | 98.12/88.19 | 97.95/81.95 | 98.22/87.82 | 98.53/89.36 | 97.60/64.07 | 76.77/28.99 | 35.56/8.48 | 99.72/91.26 | 98.47/61.83 | |||||||||||
Garden_LT_01 | 98.40/87.89 | 87.30/79.01 | 87.68/82.31 | 75.35/64.21 | 74.40/63.33 | 48.96/39.26 | 86.10/39.30 | 94.76/81.20 | 76.18/35.97 | |||||||||||
Garden_LT_02 | 98.06/87.85 | 82.98/78.21 | 86.54/79.57 | 70.89/61.41 | 70.09/64.41 | 53.23/36.09 | 70.79/33.74 | 87.30/80.75 | 73.33/33.17 | |||||||||||
Garden_LT_03 | 95.08/83.31 | 94.44/79.91 | 90.19/80.34 | 84.94/77.60 | 67.55/46.29 | 53.70/43.79 | 61.88/28.28 | 89.20/81.52 | 78.73/47.56 | |||||||||||
Garden_LT_04 | 97.03/86.69 | 94.60/86.45 | 91.64/81.17 | 85.36/73.77 | 69.24/55.00 | 40.73/39.47 | 73.55/31.12 | 92.61/82.24 | 84.80/61.57 | |||||||||||
Average | 98.03/86.99 | 94.68/82.73 | 94.11/83.14 | 90.13/81.03 | 85.60/64.82 | 70.24/46.47 | 64.71/23.98 | 94.61/87.91 | 89.85/66.47 |
1
The best result is highlighted in Blue, the second-best result is highlighted in Red, and the third-best result is highlighted in Bold.
2
For methods supervised by translation equivariance hypothesis (Hypothesis 1), such as UniLGL Dis. Sup. and BEVPlace++ Dis. Sup., point cloud pairs with distances below are chosen as positive place recognition samples. For other methods listed in Section IV-B, point cloud pairs with are chosen as positive place recognition samples.
V-A1 Ablation Study
The Top-1 recall of each method listed in Section IV-B is shown in Table II. From the results, the proposed method achieves an average Top-1 recall of and when using FoV-limited LiDAR and panoramic LiDAR, respectively. Thanks to the sensor-type uniformity supervision strategy based on the viewpoint invariance hypothesis proposed in Section II-B, the proposed UniLGL achieves consistent place recognition performance across heterogeneous LiDAR sensors. In contrast, methods trained under the conventional translation equivariance hypothesis, such as UniLGL Dis. Sup. and BEVPlace++ Dis. Sup., exhibit an approximate drop in recall and a drop in precision on sequences collected using FoV-limited LiDARs compared to those using panoramic LiDARs. It is worth noting that UniLGL Dis. Sup. outperforms UniLGL in certain sequences in terms of place recognition performance. This is primarily attributed to the difference in criteria used for determining positive place recognition samples. Specifically, as the example given in Fig. 4, UniLGL Dis. Sup. employs a distance-based criterion, which may neglect far-apart point cloud pairs that have significant overlap. To evaluate the impact of the local feature loss on LPR performance, an ablation study is also conducted by comparing models with and without viewpoint-invariant local feature supervision. From the comparison results, UniLGL consistently outperforms UniLGL w/o Loc. Feat. on all sequences collected using FoV-limited LiDARs in terms of both recall and precision. This demonstrates that the proposed viewpoint invariance supervision strategy effectively equips LPR networks with sensor-type uniformity. For panoramic LiDARs, removing local feature supervision encourages the network to focus more on learning global descriptors, which leads UniLGL w/o Loc. Feat. to outperform UniLGL on certain sequences. However, for sequences with long-term time intervals, such as Garden_LT_01, Garden_LT_02, Garden_LT_03, and Garden_LT_04, UniLGL achieves an improvement of in recall and in precision when compare with UniLGL w/o Loc. Feat., demonstrating that the viewpoint-invariant supervision of pixel-level local features effectively enhances the generalization capability of LPR. To illustrate the effectiveness of spatial and material uniformity, we compare the place recognition performance of UniLGL with UniLGL w/o intensity and UniLGL w/o spatial. As shown in Table II, UniLGL effectively fuses the spatial and intensity information of point clouds through the feature fusion network introduced in Section II-A, significantly enhancing place recognition performance on sequences collected by heterogeneous LiDARs across campus, urban driving, and garden scenarios. As shown in the Precision-Recall curve in Fig. 5, the UniLGL consistently outperforms all ablation variants.
V-A2 Benchmark with Baseline Methods
As the results summarized in Table II, benefiting from the consideration of the uniformity introduced in Section II, UniLGL delivers consistent LPR performance across all 16 sequences. Compared to SOTA learning-based[15, 37] and handcrafted[53] LGL methods, it achieves improvements of – in recall and – in average precision on FoV-limited LiDAR collected sequences, while on panoramic LiDAR collected sequences, the recall and average precision improve by – and –, respectively. We also compare the place recognition performance of the proposed UniLGL and BEVPlace++ under the conventional translation equivariance hypothesis. The results show that UniLGL Dis. Sup. consistently outperforms BEVPlace++ Dis. Sup. in both recall rate and average precision across various LiDAR configurations. As the precision-recall curves shown in Fig. 5, the proposed UniLGL consistently outperforms the other three SOTA LPR methods across all 16 data sequences. In Fig. 6, the effectiveness of UniLGL is further demonstrated across a series of challenging scenarios. The results show that UniLGL successfully retrieves correct matches in situations involving reversed loops, large viewpoint variations, loops with low overlap regions, sparse and repetitive environments, and long-term revisits, where other methods are prone to fail. The underlying reason is that explicitly embedding uniformity into the network enables the framework to learn robust place recognition using geometric and material information simultaneously across heterogeneous LiDARs.
V-B Evaluation of Complete Global Localization
In this section, experiments are conducted to evaluate the accuracy of complete global localization, which estimates the global pose of query point cloud on against database references without prior knowledge of the initial pose. For each query point cloud, UniLGL retrieves its Top-1 match from the database via place recognition and computes the global pose using the pose estimation algorithm derived in Section III. Three metrics are introduced to evaluate the global localization performance of all methods listed in Section IV-B.
Translation Error: measures the Euclidean distance between estimated and ground truth translation vectors.
(24) where and denote the estimated and ground truth translation vectors, respectively.
Rotation Error: measures the difference between the estimated and ground truth rotation.
(25) where denotes the mapping from a rotation matrix to a rotation vector, and and denote the estimated and ground truth rotation matrix, respectively.
Success Rate: represents the fraction of scan pairs with translation error and rotation error .
V-B1 Ablation Study
Sequence | UniLGL |
BEVPlace++ | Logg3D-Net | RING++ | |||||||
FoV-limited LiDAR |
Mid_NTU_02 | 83.92/1.12/0.59 | 80.47/1.21/0.60 | 80.25/1.11/0.69 | 35.44/1.49/0.71 | 55.19/1.94/0.57 | 62.13/1.08/0.38 | 31.02/2.07/0.85 | ||||||
Mid_NTU_10 | 91.06/0.66/0.35 | 90.44/0.70/0.41 | 87.77/0.67/0.45 | 80.07/0.98/0.44 | 88.70/1.10/0.31 | 93.75/0.78/0.28 | 57.08/1.45/0.78 | |||||||
Mid_NTU_13 | 80.76/1.31/0.59 | 77.57/1.37/0.62 | 78.32/1.28/0.68 | 42.91/1.29/0.52 | 55.31/1.77/0.51 | 58.15/1.33/0.41 | 32.74/1.65/0.81 | |||||||
Garden_01 | 91.44/0.63/0.28 | 90.64/0.65/0.33 | 89.17/0.63/0.32 | 80.42/0.71/0.32 | 91.59/0.68/0.18 | 91.77/0.63/0.31 | 83.33/1.12/0.67 | |||||||
Garden_02 | 91.58/0.70/0.32 | 90.59/0.70/0.35 | 91.00/0.73/0.36 | 73.62/0.86/0.38 | 89.05/0.78/0.20 | 91.47/0.76/0.34 | 83.42/1.22/0.68 | |||||||
Garden_03 | 90.55/0.71/0.31 | 86.70/0.73/0.34 | 87.15/0.72/0.37 | 69.35/0.91/0.41 | 85.35/0.79/0.20 | 87.58/0.71/0.33 | 67.12/1.24/0.73 | |||||||
Garden_04 | 95.34/0.71/0.32 | 95.16/0.77/0.35 | 88.24/0.72/0.34 | 76.87/0.83/0.35 | 90.00/0.77/0.20 | 95.79/0.71/0.33 | 74.23/1.23/0.74 | |||||||
Average | 89.24/0.83/0.39 | 87.37/0.88/0.43 | 85.99/0.84/0.46 | 65.53/1.01/0.45 | 79.31/1.12/0.31 | 82.95/0.86/0.34 | 61.28/1.43/0.75 | |||||||
Panoramic LiDAR |
OS_NTU_02 | 99.82/0.69/0.27 | 99.74/0.66/0.28 | 98.45/0.55/0.45 | 63.93/1.10/0.41 | 96.15/1.97/0.42 | 86.56/0.79/0.27 | 75.51/1.44/0.50 | ||||||
OS_NTU_10 | 99.66/0.37/0.18 | 99.65/0.38/0.19 | 98.88/0.36/0.26 | 96.37/0.51/0.22 | 99.38/1.10/0.17 | 99.07/0.43/0.16 | 75.57/1.44/0.50 | |||||||
OS_NTU_13 | 100.0/1.06/0.44 | 99.86/1.08/0.48 | 99.87/1.05/0.54 | 60.90/1.53/0.60 | 96.33/1.86/0.52 | 78.15/1.50/0.50 | 58.65/1.91/0.60 | |||||||
Snail_81R_02 | 79.81/0.83/1.04 | 78.60/0.82/1.07 | 63.30/1.03/0.98 | 11.79/1.82/1.11 | 73.06/1.27/1.06 | 31.51/2.07/0.90 | 29.88/1.40/0.94 | |||||||
Snail_81R_03 | 81.82/0.75/0.50 | 82.70/0.79/0.57 | 82.19/0.75/0.63 | 14.51/1.89/0.90 | 80.39/1.31/0.65 | 40.57/2.03/0.64 | 31.46/1.52/0.61 | |||||||
Garden_LT_01 | 81.13/1.21/0.52 | 64.24/1.36/0.57 | 56.39/1.37/0.71 | 1.69/2.20/0.75 | 13.17/1.99/0.77 | 19.75/1.80/0.52 | 77.23/1.65/0.48 | |||||||
Garden_LT_02 | 75.16/1.24/0.67 | 61.93/1.28/0.55 | 51.89/1.41/0.74 | 1.08/1.88/0.71 | 17.46/1.29/0.77 | 20.20/1.91/0.56 | 61.35/1.62/0.47 | |||||||
Garden_LT_03 | 76.28/1.30/0.53 | 71.62/1.46/0.56 | 55.40/1.63/0.68 | 4.38/2.31/0.87 | 18.07/1.23/0.75 | 26.40/1.81/0.56 | 53.21/1.58/0.44 | |||||||
Garden_LT_04 | 79.39/1.30/0.56 | 77.42/1.39/0.57 | 57.81/1.67/0.67 | 5.32/1.68/0.89 | 18.25/1.30/0.79 | 21.32/1.57/0.55 | 61.61/1.60/0.48 | |||||||
Average | 85.90/0.97/0.52 | 81.75/1.02/0.54 | 73.80/1.09/0.63 | 28.89/1.66/0.72 | 56.92/1.51/0.66 | 47.06/1.55/0.52 | 58.27/1.57/0.56 |
For the results shown in Table. III, the proposed UniLGL achieves an average success rate of and when using FoV-limited LiDAR and panoramic LiDAR, respectively. To validate the effectiveness of introducing spatial and material uniformity, we compare UniLGL with two ablated variants, UniLGL w/o Intensity and UniLGL w/o Spatial. The proposed UniLGL achieves a – and – improvement in global localization success rate compared to the respective variants that exclude material and spatial uniformity on sequences collected by FoV-limited LiDAR and panoramic LiDAR, respectively. For challenging sequences with long time intervals, such as Garden_LT_01 to Garden_LT_04, UniLGL achieves more robust global localization performance, with average improvements of and in success rate over UniLGL w/o Intensity and UniLGL w/o Spatial, respectively. For large-scale urban driving sequences such as Snail_81R_02 and Snail_81R_03, UniLGL w/o Intensity and UniLGL w/o Spatial achieve comparable global localization success rates to UniLGL on Snail_81R_03, but are outperformed by UniLGL on Snail_81R_02. This discrepancy can be attributed to the difference in dynamic object density, where Snail_81R_02 was recorded during peak traffic hours, while Snail_81R_03 was collected at night during off-peak hours. These results indicate that UniLGL offers a more robust global localization to dynamic objects, such as moving vehicles and pedestrians, compared to the ablated variants. To illustrate the effectiveness of introducing viewpoint invariance at the local feature level, an ablation study is also conducted by comparing models with and without viewpoint-invariant local feature supervision. From the comparison result, UniLGL consistently outperforms UniLGL w/o Loc. Feat. across all 16 sequences, achieving an average improvement of – in global localization success rate on both FoV-limited and panoramic LiDAR. For metric localization, UniLGL fuses pixel-level viewpoint-invariant local features from spatial and intensity BEV images and maps them to the point cloud, further delivering more accurate 6-DoF metric localization compared with its ablated variants.
V-B2 Benchmark with Baseline Methods
Table III presents the success rate, translation error, and rotation error of global localization for each method listed in Section IV-B. The proposed method performs pixel-wise matching of corresponding BEV images by conducting a KNN search in the local feature space and maps the matched pixel pairs to corresponding point pairs to estimate robust 6-DoF poses on SE(3) using the graduated non-convexity algorithm. This approach enables high-precision global localization without prior knowledge of the initial pose, improving the global localization success rate by – and – compared to SOTA BEV-based LGL methods, BEVPlace++ and RING++, on sequences collected by FoV-limited LiDAR and panoramic LiDAR, respectively. It is worth noting that Logg3D-Net outperforms the proposed UniLGL on certain sequences, as Logg3D-Net is enhanced by its follow-up work SpectralGV, which re-ranks the Top-20 retrieval candidates based on local feature consistency. In contrast, UniLGL incorporates local feature consistency directly within the network, allowing it to achieve outstanding global localization performance using only the Top-1 retrieval. As shown in Table III, for long-term sequences such as Garden_LT_01 to Garden_LT_04, the hand-crafted LGL method RING++ achieves significantly better global localization performance compared to learning-based methods BEVPlace++ and Logg3D-Net. This is mainly attributed to the inherent generalization ability of handcrafted descriptors and features. In contrast, the proposed UniLGL guided the feature fusion network introduced in Section II-A using the viewpoint invariance hypothesis (Hypothesis 2), which encodes spatial, material, and sensor-type uniformity into the LGL process. These designs enable UniLGL to achieve superior generalization performance compared to existing SOTA methods. Specifically, on the Garden_LT_01 to Garden_LT_04, UniLGL improves the global localization success rate by over compared to BEVPlace++ and Logg3D-Net, and by more than even compared to RING++. Due to its higher global localization success rate compared to other methods, UniLGL considers a broader range of challenging scenarios, such as reverse loops, global localization under large rotations and translations, low-overlap regions, and sparse and repetitive long-term environments shown in Fig. 8, which are excluded by other methods due to localization failure. Despite considering these more difficult cases, UniLGL still maintains comparable pose estimation accuracy. In Table III, only the localization error for successful global localization is tabulated. For a fair comparison, the localization error for positive place recognition, including both successful and failed global localization cases, is illustrated in Fig. 7. As the results are shown in Fig. 7 and Fig. 8, compared to SOTA LGL methods, UniLGL produces fewer global localization outliers and effectively aligns pairs of point clouds in various challenging scenarios without additional registration.
V-C Evaluation of Running Time
The average processing consumption of the proposed UniLGL and each benchmark global localization method is summarized in Table IV. For a fair comparison, we evaluate the running time of all benchmark methods with an Intel i9-13900KF CPU and an NVIDIA RTX 4080 GPU. For learning-based LGL methods, UniLGL, BEVPlace++, and Logg3D-Net, the LPR networks are deployed on the GPU, while the global localization algorithms are executed on the CPU. For handcrafted LGL methods, RING++, the LPR descriptor extraction and closed-form position estimation are deployed on the CPU, whereas the global descriptor matching and the rotation estimation are accelerated through GPU-based fast Fourier transformation. UniLGL formulates the LPR problem as an image retrieval task, enabling efficient LPR through KNN search over the global descriptor space. This approach achieves consistent place recognition time consumption across varying-scale scenarios and heterogeneous LiDAR sensors. From the results shown in Table IV, UniLGL outperforms SOTA LGL methods in LPR time consumption across all 16 sequences. Compared to SOTA learning-based LGL methods, BEVPlace++ and Logg3D-Net, UniLGL adopts a Transformer as the backbone, which can be easily accelerated using Transformers acceleration toolkits (e.g., xFormers[79]) and extracts more information-dense global descriptors, leading to a - reduction in LPR time consumption. Compared with RING++, UniLGL replaces exhaustive search with global descriptor matching, which avoids the curse of dimensionality and improves computational efficiency by over . For global localization, RING++ achieves the best real-time performance because its rotation estimation is coupled with LPR, allowing the global pose to be estimated using a closed-form position-only solution. However, this coupling also leads to a curse of dimensionality in LPR. As shown in Table IV, BEVPlace++ demonstrates more stable global localization time consumption compared to UniLGL. This is attributed to BEVPlace++ performing 3-DoF global localization using FAST keypoints[80] extracted from BEV images. In contrast, the proposed method is a fully 6-DoF global localization approach, which maps all local feature correspondences from BEV images to the point cloud and solves for the full 6-DoF global pose on SE(3). Compared to the 6-DoF global localization method, Logg3D-Net, which performs re-ranking of the Top-20 LPR candidates using SpectralGV[57], followed by RANSAC-based registration on the paired raw point clouds, the proposed UniLGL conducts registration only on the matched local feature correspondences, resulting in over reduction in computational cost. UniLGL achieves an average total time consumption of and when using FoV-limited LiDAR and panoramic LiDAR, respectively. As shown in the results from Table II to Table IV, UniLGL delivers high-performance LPR and fully 6-DoF global localization while maintaining comparable real-time capability.
Sequence | UniLGL | BEVPlace++ | Logg3D-Net | RING++ | |||||||||
PR | GL | Total | PR | GL | Total | PR | GL | Total | PR | GL | Total | ||
FoV-limited LiDAR |
Mid_NTU_02 | 6.90 | 72.15 | 79.05 | 15.27 | 54.49 | 69.76 | 41.45 | 2813.88 | 2855.33 | 351.26 | 1.38 | 352.64 |
Mid_NTU_10 | 6.81 | 82.35 | 89.16 | 15.27 | 53.90 | 69.17 | 41.19 | 2346.94 | 2388.13 | 372.46 | 1.25 | 373.71 | |
Mid_NTU_13 | 6.80 | 74.41 | 81.21 | 15.25 | 53.42 | 68.67 | 42.23 | 2774.24 | 2816.48 | 355.01 | 1.25 | 356.26 | |
Garden_01 | 7.55 | 67.51 | 75.06 | 33.94 | 50.21 | 84.15 | 41.03 | 2167.09 | 2208.12 | 1301.20 | 1.27 | 1302.47 | |
Garden_02 | 7.56 | 65.74 | 73.30 | 33.88 | 50.33 | 84.21 | 41.08 | 2339.33 | 2380.41 | 1300.79 | 1.28 | 1302.08 | |
Garden_03 | 7.58 | 67.70 | 75.28 | 34.29 | 51.27 | 85.56 | 41.32 | 2469.13 | 2510.45 | 1320.16 | 1.27 | 1321.44 | |
Garden_04 | 7.90 | 66.59 | 74.49 | 33.87 | 50.98 | 84.85 | 40.97 | 2198.04 | 2239.01 | 1305.79 | 1.28 | 1307.07 | |
Average | 7.29 | 70.72 | 78.06 | 24.12 | 52.06 | 77.67 | 41.32 | 2432.60 | 2474.09 | 751.62 | 1.29 | 753.20 | |
Panoramic LiDAR |
OS_NTU_02 | 7.56 | 179.61 | 187.17 | 15.30 | 56.24 | 71.54 | 42.50 | 2998.34 | 3040.842 | 838.60 | 1.24 | 839.84 |
OS_NTU_10 | 7.53 | 207.54 | 215.07 | 15.34 | 57.09 | 72.44 | 42.39 | 2303.79 | 2346.185 | 901.14 | 1.22 | 902.37 | |
OS_NTU_13 | 7.54 | 243.55 | 251.09 | 15.27 | 57.61 | 72.88 | 43.45 | 3202.87 | 3246.32 | 966.50 | 1.24 | 967.74 | |
Snail_81R_02 | 7.69 | 151.69 | 159.39 | 24.08 | 58.05 | 82.14 | 42.57 | 3752.56 | 3795.14 | 1109.51 | 1.24 | 1110.74 | |
Snail_81R_03 | 7.58 | 141.14 | 148.73 | 18.38 | 57.33 | 75.70 | 42.66 | 3795.48 | 3838.13 | 832.40 | 1.24 | 833.64 | |
Garden_LT_01 | 8.71 | 60.58 | 69.29 | 33.83 | 51.17 | 84.99 | 37.32 | 2372.85 | 2410.17 | 1173.54 | 1.48 | 1175.01 | |
Garden_LT_02 | 9.05 | 57.36 | 66.41 | 33.81 | 50.54 | 84.35 | 37.08 | 2214.02 | 2251.10 | 1167.90 | 2.01 | 1169.91 | |
Garden_LT_03 | 8.01 | 59.46 | 67.47 | 33.84 | 51.48 | 85.33 | 37.59 | 2204.00 | 2241.59 | 1184.55 | 1.93 | 1186.48 | |
Garden_LT_04 | 9.09 | 60.00 | 69.09 | 33.87 | 51.26 | 85.13 | 37.73 | 2140.98 | 2178.71 | 1197.18 | 1.98 | 1199.16 | |
Average | 8.06 | 110.27 | 119.95 | 23.37 | 54.44 | 79.18 | 40.28 | 2707.24 | 2748.10 | 1030.55 | 1.47 | 1032.04 |
VI Real-world Applications
VI-A Application 1: LiDAR-only Localization for Self-driving Truck in Large-scale Port Scenario
To attest to its practicality, the proposed UniLGL is integrated with a multi-LiDAR odometry (CTE-MLO[8]) to deliver high-precision and comprehensive localization and mapping for a full-size autonomous driving truck. Driving tests spanning over are conducted across three highly similar yet distinct areas within a large-scale seaport. As shown in Fig. 9(a), the truck is equipped with 7 heterogeneous LiDARs, including 1 Ouster OS1-32, 3 Robosense Helios 32, and 3 Robosense M1 solid-state LiDAR. During autonomous driving tests, CTE-MLO[8] is adopted to provide real-time state feedback to the control level of the truck by tightly coupling multi-LiDAR measurements. However, for high-level planning tasks, such as port vehicle scheduling, drift-free localization and mapping is typically required. To achieve high-precision long-term localization and mapping, we introduce UniLGL to perform loop closure detection and relative pose estimation, which is used to correct the long-term drift of multi-LiDAR odometry. A factor graph is then employed to fuse the drift-prone odometry with the loop closure constraints provided by UniLGL, enabling globally consistent and accurate localization over extended periods. The mathematical expression of factor graph optimization is shown as follows:
where denotes the trajectory of the truck, and are the -th and -th poses along the trajectory, with , and are the sets of index pairs corresponding to odometry constraints and loop closure constraints, and and represent the relative pose measurements between and obtained from CTE-MLO and UniLGL, respectively. As shown in Fig. 10, UniLGL provides a high success rate of global localization results in the highly repetitive ports environment, supplying high-quality loop closure constraints for the factor graph optimization problem defined in (26). This effectively eliminates the drift of multi-LiDAR odometry and ensures reliable long-term localization and mapping performance in real-world autonomous truck driving scenarios. The high-quality point cloud shown in Fig. 10 illustrates that the proposed method can provide robust trajectory estimation to reconstruct a dense 3D, high-precision map in a large-scale port scenario when integrated with LiDAR-only odometry.
VI-B Application 2: Multi-MAV Collaborative Exploration with Bandwidth Limitation
To further validate its practicality on lightweight power-limited platforms, UniLGL is deployed on a multi-MAV system constructed by 4 identical MAVs to enable collaborative localization during an exploration task in a field scenario. As shown in Fig. 9(b), each MAV platform is equipped with a Livox Mid 360 LiDAR, a Pixhawk4 flight controller, a low-power onboard computer (Nvidia Orin NX), and a VEWOE VMA10A mesh networking unit. During collaborative exploration, we integrate the observations from UniLGL and CTE-MLO into a factor graph to enable decentralized collaborative state estimation for multiple MAVs. As illustrated in Fig. 11, the relative pose estimations provided by UniLGL are incorporated as inter-MAV measurements, while the odometry information from CTE-MLO is utilized as intra-MAV measurements within the factor graph. As illustrated in Fig. 12(a), UniLGL provides sufficient relative localization constraints among multiple MAVs, enabling simultaneous infrastructure-free collaborative localization and mapping. Notably, the use of UniLGL for inter-MAV constraints significantly reduces the communication bandwidth required in multi-agent systems. This efficiency is reflected in two main aspects: 1) UniLGL enables collaborative localization through an event-triggered mechanism, where a UAV transmits relative localization information to others only after a successful global descriptor matching, rather than continuously broadcasting point clouds as required by methods such as [81, 82, 83]. 2) For relative pose estimation, UniLGL requires transmitting only a small number of point cloud keypoints associated with matched local features, thereby eliminating the need to transmit raw point clouds for pose refinement. As shown in Fig. 12(b), the UniLGL-based collaborative localization approach achieves substantially lower communication bandwidth consumption compared to methods that lack event-triggered mechanisms (noted as Point Cloud in Fig. 12(b)) or require additional point cloud registration (noted as Point Cloud. Event in Fig. 12(b)). Owing to the superior efficiency of UniLGL in terms of computational and communication resource consumption, we integrate UniLGL-based collaborative localization with a decentralized exploration algorithm DPPM[84] and the MAV trajectory tracking control algorithm FxTDO-MPC[85], which enables a real-world exploration fully onboard. As shown in Fig. 12(c), the multi-MAV system completed high-precision scanning in a dense forest environment. The area of the exploration regions reaches .
VII Conclusion
In this article, a Uniform LiDAR-based Global Localization system, UniLGL, is developed to achieve cascaded place recognition and global pose estimation with the consideration of spatial and material uniformity as while as sensor-type uniformity. To equip UniLGL with spatial and material uniformity, we represent the 4D point cloud information in a lossless manner using an image pair consisting of a spatial BEV image and an intensity BEV image, and design an end-to-end BEV fusion network for place recognition. For sensor-type uniformity, a viewpoint invariance hypothesis is introduced to replace the conventional translation equivariance assumption commonly used in existing LPR networks[24, 15, 14, 54, 25], which hypothesis guides UniLGL to learn global descriptors and local features with consistency across geographically distant but co-visible areas (as demonstrated in Fig. 4). Moreover, a vision foundation model, DINO[68], is elegantly integrated into the proposed BEV fusion network to enhance the generalization capability of global descriptors and local features, without requiring large amounts of LiDAR training data. Thanks to the consistency of local features across co-visible areas, a global pose estimator is derived using graduated non-convexity optimization[74] to estimate the 6-DoF global pose based on point-level correspondences established through local feature matching between BEV images. With the successful integration of a series of theory and practical implementations (e.g. uniform LPR, task agnostic vision foundation model, and robust global estimator on SE(3)) into the UniLGL, the proposed method is demonstratively competitive compared to SOTA LPR/LGL methods. The extensive experimental results have demonstrated that the proposed UniLGL remains robust under extremely challenging conditions, such as long-term global localization and large viewpoint variations, across heterogeneous LiDAR configurations. Furthermore, the proposed UniLGL has been deployed to support autonomy on diverse platforms, from full-size trucks to power-limited MAV, which enables high-precision truck localization and mapping in a port environment and multi-MAV collaborative exploration in a forest environment. These real-world deployments affirm the extendability and applicability of UniLGL in industrial and field scenarios.
Appendix A Effectiveness of introduce Foundation Model into LGL & Zero-Shot Generalization across FoV-Limited and Panoramic LiDAR
Foundation models in robotics aim to provide systems with broad generalization capabilities by leveraging large-scale pretraining on diverse data sources. To facilitate understanding of the role of introducing foundation models into LGL, we evaluate the LGL performance of three variants: UniLGL, UniLGL initialized with a foundation model but without fine-tuning (UniLGL w/o FT), and UniLGL trained from scratch without foundation model initialization (UniLGL w/o FM). It is worth noting that, to demonstrate the zero-shot generalization ability, no cross-modal training is performed. We perform benchmark experiments on two datasets, MCD[76] and Garden[78], both of which contain paired FoV-limited and panoramic LiDAR measurements. During the experiments, the FoV-limited LiDAR scans are used as queries, while the panoramic LiDAR scans serve as the database, allowing us to assess the cross-modal generalization ability of LGL methods. The associated data sequences are referred to as NTU_CM_XX and Garden_CM_XX, respectively.
Sequence | UniLGL |
NTU_CM_02 | 97.75/88.63 | 37.95/16.38 | 0.60/4.54 | ||
NTU_CM_10 | 92.95/83.45 | 32.61/14.76 | 3.64/5.18 | ||
NTU_CM_13 | 96.46/88.49 | 41.02/17.50 | 8.59/7.19 | ||
Garden_CM_01 | 93.19/80.62 | 26.02/15.08 | 3.33/14.84 | ||
Garden_CM_02 | 91.86/84.73 | 30.54/19.04 | 17.54/18.97 | ||
Garden_CM_03 | 94.37/81.45 | 31.32/16.65 | 19.71/15.52 | ||
Garden_CM_04 | 87.37/86.03 | 37.06/19.82 | 24.86/15.59 | ||
Average | 93.42/85.20 | 33.78/17.03 | 11.18/11.69 |
1
The best result is highlighted in Bold.
Sequence | UniLGL |
NTU_CM_02 | 86.75/0.76/1.47 | 0.07/0.78/2.34 | 0.01/1.92/3.45 | ||
NTU_CM_10 | 78.19/0.65/1.34 | 0.13/1.46/3.44 | 0.01/1.54/3.92 | ||
NTU_CM_13 | 86.13/0.69/1.49 | 0.35/1.34/3.57 | - | ||
Garden_CM_01 | 75.77/0.86/2.44 | 0.25/1.44/3.54 | - | ||
Garden_CM_02 | 76.10/0.75/2.44 | 0.34/1.22/3.27 | 0.01/1.77/4.01 | ||
Garden_CM_03 | 76.07/0.77/2.53 | 0.41/1.19/2.58 | - | ||
Garden_CM_04 | 76.94/0.67/2.37 | 0.36/1.08/2.44 | 0.01/1.27/2.89 | ||
Average | 79.42/0.74/2.01 | 0.27/1.21/3.03 | 0.01/1.63/3.57 |
1
The best result is highlighted in Bold.
2
- denotes no successful global localization achieved.
As shown in the t-SNE[61] visualization in Fig. 13, introducing a foundation model imparts a certain level of generalization ability to the LPR network, enabling UniLGL w/o FT to achieve better clustering performance than UniLGL w/o FM, without fine-tuning. By initializing the network weights with DINO[68] and fine-tuning with only a small amount of homogeneous LiDAR data, UniLGL learns highly discriminative global descriptors in the heterogeneous LPR task. In addition to the above qualitative analysis, we present quantitative results of place recognition and global localization performance in Table V and Table VI, respectively. The results show that, thanks to the introduction of the foundation model, UniLGL achieves outstanding zero-shot cross-modal generalization ability. When compared to UniLGL w/o FT and UniLGL w/o FM, UniLGL achieves a - improvement in recall and a - improvement in average precision. For global localization, UniLGL achieves over increase in successful localization rate compared to UniLGL w/o FT and UniLGL w/o FM, with average global localization accuracy reaching and .